Systems Staff Meeting Notes
July 31, 2006
July 31, 2006
Present: Cynthia Bowker, Will Chaney, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Geoffrey Stewart, Ying Tang, Stephen WilliamsAnnouncements
Ongoing Systems Updates
- Cynthia cleaned her carpets this weekend. Bobbie cleaned the mold from the track of her shower stall this weekend.
- Systems construction - phone number for the crew supervisor, Barry: 678-431-1282. Need to give him a call before going down to the area. Rob got a quote on the mold removal/painting.
- Martha's last day is tomorrow. Beach party, 3rd floor East.
- Library West opening 1PM Wed Aug 2. Construction on Starbucks to begin soon. Room 1 to close as soon as West is open.
- Order for four DVD/VCR combos for West - they are going to be back ordered, so we will switch to a different, available brand. Same vendor (Dell).
- Need port list for all of the Libraries to verify charging for wall plate service. We've received them from Ed & MSL - nothing for East.
- ALEPH access given to a student with no networking sign-on - just need to remind liaisons as to the security breach of signing people on to the network under someone else's sign-on.
Software Projects Status
- Linux agent downloaded, but has already been superseded. Backup Exec requires a newer copy -43 mgs of Linux formatted files.
- West high-end computers - very slow in performing. Some tweaks have been found. RAM cache on drives need to be equal. Photoshop scratch file needs to be on another drive. Need high-end graphics card.
- Upcoming - West opening - Systems should be available for anything.
- Media Tech - Aug 11-30 - they will be setting up the West classrooms/training room. They will work over the weekend. Facilities will give them access.
- Quality Control document is in the works.
- Grant tracking - still under review.
- Training database - getting feedback from HR & HSSS.
- ERES - meeting tomorrow with Michelle Foss. Last meeting before the roll-out in the fall.
- Personnel database revamping - Winston is working on it.
- Greenstone - Ying is doing some last-minute fixing before she leaves.
- Mark is on vacation this week.
Students and Departments
- Circ manual is done.
- Contact for MetaLib is Carol Turner. Waiting for CM to come back with a list of approved subjects. It will stay in beta testing for the fall.
- Test blogging server to be set up in Tom's office.
- All three students should be returning for the fall. No schedule as of yet.